OpenStudy (anonymous):

please check my work ^.^ medal + fan :] if i got it wrong don't be too harsh, maths is not my best subject

OpenStudy (anonymous):

@campbell_st @phi

OpenStudy (anonymous):

I dont know if this is right or not, but please go back and recheck what the problem is saying, i think you may have made a mistake.

OpenStudy (anonymous):

Well are you 100% sure about that? Cause lol I'm not just going to go back on an assumption. No offense but if you are not sure then I won't redo my work.

OpenStudy (anonymous):

Im 100% sure :)

OpenStudy (campbell_st):

your solution is correct

OpenStudy (anonymous):

Oh wow really @campbell_st ? o.O you sure? lol I trust ya but like I seriously can't believe I did something right?

OpenStudy (anonymous):

Did you check it? becuase sometimes my nname deceives me, and other times im just plane wong, so i would listen to Campbell . Sorry!

OpenStudy (campbell_st):

the perpendicular distance from the directrix to the focus is 6 units so the vertex is halfway between the focus and directrix on the line x = -5 so the vertex is at (-5, 2) so half od 6 is 3 this is the focal length, a = 3 the version of the equation I use is \[(x - h)^2 = 4a(y - k)\] where (h, k) is the vertex and a is the focal length so substittuting you get \[(x + 5)^2 = 4 \times 3(y - 2)\] now make y the subject hope it helps

OpenStudy (mathstudent55):

You are correct.

OpenStudy (anonymous):

Thanks for that info, it does help! :]

OpenStudy (anonymous):

@mathstudent55 thank you :]

OpenStudy (anonymous):

I can't give both of you medals :[

OpenStudy (mathstudent55):

\(\sqrt{(x + 5)^2 + (y -5)^2} = \sqrt{(y + 1)^2} \) \((x + 5)^2 + (y -5)^2 = (y + 1)^2 \) \((x + 5)^2 + \cancel{y^2} -10y + 25 = \cancel{y^2} + 2y + 1\) \((x+5)^2 + 24 = 12y\) \( 12y =(x+5)^2 + 24 \) \(y = \dfrac{1}{12}(x + 5)^2 + 2\)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

Where did the problem go? Keep the problem the way it was because this may help other people.